Velocity of a particle is expressed as
Vx=350-200t²m/s
(a)=find average acceleration
(b)=instantaneous acceleration
1
Expert's answer
2020-03-18T10:17:49-0400

(a) The average acceleration


"a_x=\frac{v_2-v_1}{t_2-t_1}=\frac{(350-200t_2^2)-(350-200t_1^2)}{t_2-t_1}""=-200(t_2+t_1)."


(b) The instantaneous acceleration


"a_x(t)=v_x'(t)=(350-200t^2)'=-400t."